Spt X Company is considering replacing one of its machines in order to save operating costs. Operating costs with the currentFor 11, i chosed 3, is it correct?
Answer #1

Net Present value = Present value of cash inflows – Present value of cash outflows

= (-65000+6000)+ (60,000-48000)*PVAF(7%, 6 years) + 1,000*PVF(7%, 6 years)

= - 59000 + 12000*4.767 +1000*0.666

= -$1,130

i.e. B

NPV = 159000*PVF(6%, 1 year) + 159000*PVF(6%, 2 years) + 112,000*PVF(6%, 3 years) + 112,000*PVF(6%, 4 years) – 89000 – 370000

= 159000*0.943 + 159000*0.890 + 112000*0.840 + 112000*0.792 – 89000 – 370000

= $15,791

i.e. A approx.

Extra cost of Machine B = 58,000-50,000 = $8000
Year Incremental Cash flows Cumulative Cash flows
0 -8000 -8000
1 -1000 -9000
2 4000 -5000
3 5000 0
4 5000 5000
5 3000 8000
6 3000 11000
7 2000 13000
Hence, payback period = 3 years
i.e. B
Payback period is the numbers of years taken to recover extra cost